Solve for X, Y, and Z in the following system of three equations:
X + Y + 2Z = 8
X + Y = 4
X + 3Y + Z = 8

Answer by checkley71(8403)

X+Y=4 THEN 4+2Z=8 OR 2Z=4 OR Z=2 THEN X+3Y+2=8 OR
X+3Y=8
X+Y=4 SUBTRACTING WE GET
-------
2Y=4 OR Y=2 THE X+2=4 OR X=2
IT'S NOT AN EASY WALK-THROUGH. EVERY SET OF EQUTIONS TAKES A DIFFERENT APPROACH. SOME ARE OBVIOUS BUT OTHERS TAKE A LITTLE MORE SKILL AND IT IS SOMTHING YOU HAVE TO LEARN BY SOLVING MANY DIFFERENT TYPES. THE SECRET IS TO ELIMINATE ONE UNKNOWN AND THEN EQUATE ONE OF THE RAMAINING TERMS TO THR THIRD UNKNOWN AND SUBSTITUTE IN THE UNUSED EQUATION.